(15 pts) You have borrowed $100,000 at an annual interest rate of 4%, over 4 years, paying back the loan on a monthly basis.   At what payment (e.g., payment 5, payment 6, etc.) are you paying close to $275 in interest for that period only?   This requires setting up the function and doing some “what if” analysis by changing values / different scenarios.  

Please show your work. Thanks
Answer #1

Monthly Interest Rate = 0.04/12 = 0.0033

Paying interest $275 comes when,

Future Value = 275/0.0033 = $82,500

So,

First Calculating Monthly Payment,

PMT = [FV = 0, PV = 100,000, T = 48, I = 0.04/12]

PMT = $2,257.91

Now we need to find Time When Future Value = $82,500

so,

T = [PV = 100,000, FV = 82,500, PMT = -2,257.91, I = 0.04/12]

T = 9

So,

On 10th Payment on will pay $275 interest.